# Black Hole Superradiance of Interacting Multi-Field

### v1

@@ I Introduction
Bosonic fields around a rapidly rotating black hole manifest an instability known as black hole superradiance [1, 2, 3, 4, 5]. In particular, ultralight bosons can extract angular momentum and energy from an astrophysical black hole, forming a massive cloud around it. With recent advances in gravitational wave detection and black hole imaging, black hole superradiance has emerged as a powerful probe of the dark sector in particle physics [6, 7]. Extensive studies have explored the observational consequences of black hole superradiance, including black hole spin down [8], gravitational wave emission from the clouds [9, 10, 11], and dynamical signatures of clouds in black hole binaries [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], leading to stringent constraints on ultralight bosonic fields [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@ II Superradiance of free multi-fields
In this section, we investigate the superradiance process of non-interacting multi-fields. For the purpose of demonstration, we shall start with a model of two massive scalar fields in the Kerr background, the Lagrangian of which is given by
@@ II.1 Eigenstates of free fields
$\psi\left(x^{\mu}\right)=e^{-i\omega t}R_{\ell m}(r)S_{\ell m}(\theta)e^{im\varphi}\,$, (2)
where $x^{\mu}=(t,r,\theta,\varphi)$ are the Boyer-Lindquist coordinates, $R_{\ell m}(r)$ is the radial function, and $S_{\ell m}(\theta)$ turns out to be the spheroidal harmonics. The boundary conditions at the black hole horizon and at spatial infinity single out a set of eigenfrequencies. For bound states, i.e., states vanishes at spatial infinity, the eigenfrequencies $\omega$ are shown to be a set of discrete complex number $\omega_{n\ell m}$, labelled by three quantum numbers $n$, $\ell$ and $m$. These eigenfrequencies can be obtained numerically. For instance, Fig. 1 shows the eigenfrequencies of some bound states obtained with the continued fraction method [51]. For $\alpha\ll 1$, we have [14]
@@
$\left.+\frac{2a_{*}m\alpha^{5}}{n^{3}\ell(\ell+1/2)(\ell+1)}+{\cal O}(\alpha^{6})\right)\,$,
@@
$\times\prod^{\ell}_{j=1}\left(j^{2}(1-a^{2})+(ma-2\tilde{r}_{+}\alpha)^{2}\right)\,$,
where $a$ is the black hole dimensionless spin, $\tilde{r}_{+}\equiv 1+\sqrt{1-a^{2}}$, and $\Omega_{H}\equiv a/2M\tilde{r}_{+}$ is the angular velocity of the outer horizon. [^1] Also see Refs. [54, 55] for the improved analytical superradiance solutions. Therefore, when $m\Omega_{H}>\omega_{n\ell m}$, the bound state experiences the superradiant growth with a rate of $\Gamma_{n\ell m}\equiv{\rm Im}\,\omega_{n\ell m}$.
@@
${\bar{\psi}}_{n\ell m}=\sqrt{N_{n\ell m}}\,u_{n\ell m}({\bf r})e^{-i\delta_{n\ell m}t}\,$, (8)
where $\delta_{n\ell m}\equiv\omega_{n\ell m}-\mu$, $N_{n\ell m}$ is the occupation number of particle in the eigenstate, and $u_{n\ell m}({\bf r})$ is normalized to 1. To the leading order in $\alpha$, $u_{n\ell m}({\bf r})$ is given by the normalized hydrogenic eigen-wavefunction with a Bohr radius of $r_{B}\equiv GM/\alpha^{2}$.
@@
$\psi_{n\ell m}\propto u_{k\ell m}({\bf r})e^{-i\omega_{k\ell m}t}\,.$ (9)
The above discussion can be easily extend to the $\varphi$ field by replacing $\alpha$ with $q\alpha$.
Figure 2: Parameter space of two non-interacting fields. Without loss of generality, we assume $\psi_{211}$ to be the fastest growing mode, which excludes the white regime. Representative examples of the evolution in different colored regions are shown in Fig. 3. In particular, we find that superradiant growth of $\psi_{211}$ might be affect by $\varphi$, as which may accelerate black hole spin, causing $\psi_{211}$ depletes prematurely.
Figure 3: Presentative examples of non-interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 2, while the lower panel shows the evolution of black hole spin.
@@ II.2 Superradiance Evolution
The superradiant growth of the fields back reacts on the background geometry, leading to black hole spin down. Under the adiabatic approximation, the spin and mass of the black hole evolve as
$\dot{a}=-\sum_{i}m_{i}\gamma_{i}\epsilon_{i}$ (10)
$\frac{\dot{M}}{\mu}=-GM^{2}\sum_{i}\gamma_{i}\epsilon_{i}\,$, (11)
@@
$\gamma_{211}=\frac{1}{24}\alpha^{8}\left[a-2\alpha\left(1+\sqrt{1-a^{2}}\right)\right]\,$, (12)
and first dominates the black hole spin evolution. It continues growing until the black hole spin drops below $a_{\rm crit}=4\alpha/(1+4\alpha^{2})$, namely the superradiant condition is no longer satisfied. At this time, $\psi_{211}$ saturates with an occupation number of $\epsilon_{\rm max}=(a_{0}-a_{\rm crit})\sim 1$.
In the case of two fields, the superradiance evolution depends on the parameters. In Fig. 2, we show the parameter space of different types of evolution, assuming $\psi_{211}$ to be the fastest growing mode, [^2] The case that $\varphi_{211}$ grows faster can be obtained by replacing $q$ with $1/q$. while a representative example of each type is shown in Fig. 3. These examples are obtained by solving Eq. (10) together with $\dot{\epsilon}_{i}=\gamma_{i}\epsilon_{i}$, assuming the initial conditions of $\epsilon_{i}=10^{-10}$ when the black hole spin $a_{0}=0.99$. For simplicity, we treated the black hole mass (and hence $\alpha$) to be constant, given the fact that the black hole mass typically varies by a small amount during superradiant evolution. We find that, in region I and region IV, $\psi_{211}$ finally depletes after reaching the maximum occupation number, due to the development of $\varphi$ field.
@@ III Cubic Interactions
Assuming both fields are initially in vacuum states with small quantum fluctuations, the superradiance process can be investigated perturbatively: At the very beginning, interactions between the field is suppressed due to the low occupation numbers, and the superradiance process is the same as it in the case of free field. As superradiance continues, the occupation number in the fastest growing mode, i.e., ${\bar{\psi}}_{211}$, increases, resulting in a notable interaction between the two fields. Depending on the parameters, the interaction may lead to different processes. The effects on the superradiant growth of such processes can be investigated by their corrections on the eigenfrequncies, which will be discussed as follows.
We shall take the non-relativistic ansatz (6) as in the case of free field, and work in the weak field limit. Then Eqs. (13) and (14) can be written as
$\left(i\partial_{t}+\hat{H}\right)\left({\bar{\psi}}e^{-i\mu t}+{\bar{\psi}}^{*}e^{i\mu t}\right)=-2\bar{\lambda}\left({\bar{\psi}}e^{-i\mu t}+{\bar{\psi}}^{*}e^{i\mu t}\right)\varphi$ (15)
$\left(\partial_{t}^{2}-\hat{M}-\nu^{2}\right)\varphi=\bar{\lambda}\left({\bar{\psi}}^{2}e^{-2i\mu t}+2{\bar{\psi}}{\bar{\psi}}^{*}+{{\bar{\psi}}}^{*2}e^{2i\mu t}\right)\,$,
where $\bar{\lambda}\equiv\lambda/2\mu$ is the reduced coupling strength, and $\hat{M}\approx-\nabla^{2}-(2q\alpha\nu)/r$ to the first order in $GM/r$. Although working in the weak field limit, it is useful to consider $\hat{H}$ and $\hat{M}$ as non-Hermintian operators in order to take into account the energy exchanges between the fields and the black hole that do not manifest in the weak field limit.
To compute the corrections on the eigenfrequency, we take one of the $\bar{\psi}$} in the source terms as background, i.e., ${\bar{\psi}}_{0}=\sqrt{N_{0}}u_{0}(\mathbf{r})e^{-i\delta_{0}t}$. While ${\bar{\psi}}_{0}$ is in the ${\bar{\psi}}_{211}$ mode which should in principle grow with time, we shall consider ${\bar{\psi}}_{0}$ to be stationary in our perturbative calculation, and take $\delta_{0}$ to be the real part of $\omega_{211}-\mu$. It is convenient to use Feynman-like diagrams to represent the channels involved in Eqs. (LABEL:eq:eomm). In particular, $\bar{\psi}$} or $\bar{\varphi}$} on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the l.h.s. of the vertex, and vice versa, while $\bar{\psi}$$^{*}$ or $\bar{\varphi}$$^{*}$ on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the r.h.s. of the vertex, and vice versa. Here $\bar{\varphi}$} is defined as $\varphi={\bar{\varphi}}+{\bar{\varphi}}^{*}$. Focusing on interaction corrections on the fastest growing mode ${\bar{\psi}}_{211}$, Eqs. (LABEL:eq:eomm) involve three possible channels, which are shown in Fig. 4.
@@
$\psi_{211}$$\psi_{211}$$\psi_{211}$$\psi_{\infty}$
Figure 4: Possible channels involved in Eqs. (LABEL:eq:eomm). From left to right, the diagrams correspond to the s-channel, the t/u channel and the $3{\bar{\psi}}$ channel. As we are considering self-interaction correction, we do not distinguish t and u channel in the diagram.
@@
Taking the s-channel for instance, we have
$\left(i\partial_{t}+\hat{H}\right){\bar{\psi}}=-2\bar{\lambda}{\bar{\psi}}_{0}^{*}{\bar{\varphi}}e^{2i\mu t}$ (16)
$\left(\partial_{t}^{2}-\hat{M}-\nu^{2}\right){\bar{\varphi}}=\bar{\lambda}{\bar{\psi}}{\bar{\psi}}_{0}e^{-2i\mu t}\,$,
to solve which, we can take the ansatz
${\bar{\psi}}(t,\mathbf{r})$ $=\Psi(\mathbf{r})e^{-i(\delta_{0}+\delta\omega_{211})t}$ (17)
$\bar{\varphi}(t,\mathbf{r})$ $=\Phi(\mathbf{r})e^{-i\sigma t}\,.$
To the zeroth order of $\bar{\lambda}$, we have $\Psi=u_{0}$ and $\Phi=0$. To the first order of $\bar{\lambda}$, we have $\sigma=2\omega_{0}+\delta\omega_{211}$ and
$\left(\epsilon_{0}+\delta\omega_{211}+\hat{H}\right)\Psi=-2\bar{\lambda}\sqrt{N_{0}}u_{0}^{*}\Phi$ (18)
$\left(-\sigma^{2}-\hat{M}-\nu^{2}\right)\Phi=\bar{\lambda}\sqrt{N_{0}}u_{0}\Psi\,.$
In order to solve $\Phi$, we further wirte
$\Phi=\sum_{n,\ell,m}^{\infty}b_{n\ell m}v_{n\ell m}+\sum_{\ell,m}\int_{0}^{\infty}b_{\ell m}(k)v_{k\ell m}\mathrm{d}k\,$, (19)
where $v_{n\ell m}$ and $v_{k\ell m}$ are the eigenfunctions satisfying
$\hat{M}v_{n\ell m}=\kappa_{n}^{2}v_{n\ell m},\quad\hat{M}v_{k\ell m}=-\kappa^{2}v_{k\ell m}$ (20)
with $\kappa_{n}\equiv\sqrt{\nu^{2}-\sigma_{n}^{2}}$ and $\kappa\equiv\sqrt{\sigma^{2}-\nu^{2}}$ being the eigenvalues. Then $\Phi$ in Eqs. (LABEL:eq:sschannel) can be solved with the orthogonal conditions. Specifically, integral over the angular part indicates the nontrivial contributions only come from the modes with $\ell=m=2$, and their corresponding coefficients are
$b_{n}$ $=\frac{\bar{\lambda}\sqrt{N_{0}}}{-(2\omega_{0})^{2}-\kappa_{n}^{2}+\nu^{2}}\int u_{0}^{2}v_{n22}^{*}\mathrm{d}^{3}\mathbf{r}$ (21)
$b(k)$ $=\frac{1}{2\pi}\frac{\bar{\lambda}\sqrt{N_{0}}}{-(2\omega_{0})^{2}+\kappa^{2}+\nu^{2}}\int u_{0}^{2}v_{k22}^{*}\mathrm{d}\,,^{3}\mathbf{r}$
where we have omitted the subscriptions $\ell m$} for short. Then by multiplying $u_{0}^{*}$ on the equation of $\Psi$ and integrating over $\mathrm{d}^{3}\mathbf{r}$ on both sides, we find $\Psi=u_{0}$ to the leading order of $\bar{\lambda}$. Also, the frequency correction from the s-channel can be written as $\delta\omega_{211}^{s}=\delta\omega_{211}^{sB}+\delta\omega_{211}^{sE}$ with
@@
${\rm Im}\delta\omega_{211}^{sB}\simeq\alpha^{3}\bar{\lambda}^{2}N_{0}\frac{4q^{2}\mu}{(q^{2}-4)^{2}}\sum_{n\geq 3}\frac{{\rm Im}\sigma_{n}}{\nu}{\rm I}_{n}^{B}(q)\,$, (23)
which indicates that the superradiant growth rate of ${\bar{\psi}}_{211}$ is suppressed if it interacts with a decaying ${\bar{\varphi}}_{n22}$, as ${\bar{\varphi}}_{n22}$ may lose energy to the black hole in the s-channel, and vice versa. Here
@@
is a dimensionless factor depending only on the mass ratio $q$, with its values shown in Fig. 5. For the corrections from unbound states, we find that $\delta\omega_{211}^{sE}$ has a non-zero imaginary part only if $k_{0}$ is real due to the optical theorem. Specifically, one has
${\rm Im}\delta\omega_{211}^{sE}\simeq\left\{\begin{array}[]{ll}-\bar{\lambda}^{2}N_{0}\frac{\mu\alpha^{2}}{2\sqrt{4-q^{2}}}{\rm I}^{E}_{k_{0}}(q,\alpha),&2\omega_{0}>\nu\\ 0,&2\omega_{0}<\nu\end{array}\right$. (25)

### v2

@@ I Introduction
Bosonic fields around a rapidly rotating black hole manifest an instability known as black hole superradiance [1, 2, 3, 4, 5]. In particular, ultralight bosons can extract angular momentum and energy from an astrophysical black hole, forming a massive cloud around it. With recent advances in gravitational wave detection and black hole imaging, black hole superradiance has emerged as a powerful probe of the dark sector in particle physics [6, 7]. Extensive studies have explored the observational consequences of black hole superradiance, including black hole spin down [8], gravitational wave emission from the clouds [9, 10, 11], and dynamical signatures of clouds in black hole binaries [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], leading to stringent constraints on ultralight bosonic fields [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47].
@@ II Superradiance of free multi-fields
In this section, we investigate the superradiance process of non-interacting multi-fields (also see Ref. [54] for a similar study). For the purpose of demonstration, we shall start with a model of two massive scalar fields in the Kerr background, the Lagrangian of which is given by
@@ II.1 Eigenstates of free fields
$\psi\left(x^{\mu}\right)=e^{-i\omega t}R_{\omega\ell m}(r)S_{\omega\ell m}(\theta)e^{im\varphi}\,$, (2)
where $x^{\mu}=(t,r,\theta,\varphi)$ are the Boyer-Lindquist coordinates, $R_{\omega\ell m}(r)$ is the radial function, and $S_{\omega\ell m}(\theta)$ turns out to be the spheroidal harmonics. The boundary conditions at the black hole horizon and at spatial infinity single out a set of eigenfrequencies. For bound states, i.e., states vanishes at spatial infinity, the eigenfrequencies $\omega$ are shown to be a set of discrete complex number $\omega_{n\ell m}$, labelled by three quantum numbers $n$, $\ell$ and $m$. These eigenfrequencies can be obtained numerically. For instance, Fig. 1 shows the eigenfrequencies of some bound states obtained with the continued fraction method [55]. For $\alpha\ll 1$, we have [14]
@@
$\left.+\frac{2\tilde{a}m\alpha^{5}}{n^{3}\ell(\ell+1/2)(\ell+1)}+{\cal O}(\alpha^{6})\right)\,$,
@@
$\times\prod^{\ell}_{j=1}\left(j^{2}(1-\tilde{a}^{2})+(m\tilde{a}-2\tilde{r}_{+}\alpha)^{2}\right)\,$,
where $\tilde{a}$ is the black hole dimensionless spin, $\tilde{r}_{+}\equiv 1+\sqrt{1-\tilde{a}^{2}}$, and $\Omega_{H}\equiv\tilde{a}/2M\tilde{r}_{+}$ is the angular velocity of the outer horizon. [^1] Also see Refs. [58, 59] for the improved analytical superradiance solutions, and Refs. [60, 61] for comparison with Hawking radiation. Therefore, when $m\Omega_{H}>\omega_{n\ell m}$, the bound state experiences the superradiant growth with a rate of $\Gamma_{n\ell m}\equiv{\rm Im}\,\omega_{n\ell m}$.
@@
${\bar{\psi}}_{n\ell m}=\sqrt{N_{n\ell m}}\,u_{n\ell m}({\bf r})e^{-i\xi_{n\ell m}t}\,$, (8)
where $\xi_{n\ell m}=\omega_{n\ell m}-\mu$, $N_{n\ell m}$ is the occupation number of particle in the eigenstate, and $u_{n\ell m}({\bf r})$ is normalized to 1. To the leading order in $\alpha$, $u_{n\ell m}({\bf r})$ is given by the normalized hydrogenic eigen-wavefunction with a Bohr radius of $r_{B}\equiv GM/\alpha^{2}$.
@@
$\psi_{k\ell m}\propto u_{k\ell m}({\bf r})e^{-i\omega_{k\ell m}t}\,.$ (9)
The exact expression of eigenfunctions in Newtonian potential are discussed in Appendix. A. The above discussion can be easily extend to the $\varphi$ field by replacing $\alpha$ with $q\alpha$.
Figure 2: Parameter space of two non-interacting fields. Without loss of generality, we assume $\psi_{211}$ to be the fastest growing mode, which excludes the white regime. Representative examples of the evolution in some regions are shown in Fig. 3, while the boundaries and the superradiant evolution of each region are described in Sec. II.2. In particular, we find that superradiant growth of $\psi_{211}$ might be affected by $\varphi$, as which may accelerate black hole spin down, causing $\psi_{211}$ depletes prematurely.
Figure 3: Presentative examples of non-interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 2, while the lower panel shows the evolution of black hole spin. In the left example, $\varphi_{211}$ and $\varphi_{322}$ cannot grow sufficiently become the black hole spin drops quickly due to the superradiant growth of $\psi_{211}$. In the middle example, $\psi_{211}$, $\varphi_{211}$, $\psi_{322}$ and $\varphi_{322}$ undergo efficient superradiant growth one after another and dominate the black hole spin in sequence. In the right example, the superradiant growth of $\varphi_{211}$ is suppressed by that of $\psi_{211}$, while $\varphi_{322}$ and $\psi_{322}$ can still grow in order after $\psi_{211}$ saturates.
@@ II.2 Superradiance Evolution
The superradiant growth of the fields back reacts on the background geometry, leading to black hole spin down. Under the adiabatic approximation, the angular momentum and mass of the black hole evolve as
$\dot{J}=-GM^{2}\sum_{i}m_{i}\gamma_{i}\epsilon_{i}$ (10)
$\dot{M}=-GM^{2}\mu\sum_{i}\gamma_{i}\epsilon_{i}\,$, (11)
@@
$\gamma_{211}=\frac{1}{24}\alpha^{8}\left[\tilde{a}-2\alpha\left(1+\sqrt{1-\tilde{a}^{2}}\right)\right]\,$, (12)
and first dominates the black hole spin evolution. It continues growing until the black hole spin drops below $\tilde{a}_{\rm crit}=4\alpha/(1+4\alpha^{2})$ (for $\alpha<0.5$), namely the superradiant condition is no longer satisfied. At this time, $\psi_{211}$ saturates with an occupation number of $\epsilon_{max}=\tilde{a}_{0}-\tilde{a}_{\rm crit}\sim 1$. Similarly, $\psi_{322}$ can also undergo superradiance growth until the black hole drops below $\tilde{a}^{\psi_{322}}_{\rm crit}=2\alpha/(1+\alpha^{2})$.
In the case of two fields, the superradiance evolution depends on the parameters. In Fig. 2, we show the parameter space of different types of evolution, assuming $\psi_{211}$ to be the fastest growing mode. [^2] The case that $\varphi_{211}$ grows faster can be obtained by replacing $q$ with $1/q$. Specially, if $\gamma_{\psi_{322}}>\gamma_{\varphi_{211}}$ and $\gamma_{\psi_{322}}>\gamma_{\varphi_{322}}$ (region I.A), $\psi_{322}$ will undergo efficient superradiant growth after $\psi_{211}$ saturates and dominate black hole spin down. This process is then followed by the superradiant growth of $\varphi_{211}$ and $\varphi_{322}$. If $\gamma_{\varphi_{211}}<\gamma_{\psi_{322}}<\gamma_{\varphi_{322}}$ and $\tilde{a}^{\varphi_{322}}_{\rm crit}>\tilde{a}^{\psi_{211}}_{\rm crit}$ (region I.B), $\varphi_{211}$ and $\varphi_{322}$ cannot grow sufficiently as the black hole spin drops below their critical spin quickly as $\psi_{211}$ grows, while $\psi_{322}$ can still grow sufficiently after $\psi_{211}$ saturates. If $\gamma_{\varphi_{211}}>\gamma_{\psi_{322}},\gamma_{\varphi_{322}}$ and $\tilde{a}^{\varphi_{211}}_{\rm crit}<\tilde{a}^{\psi_{211}}_{\rm crit}$ (region II), $\varphi_{211}$ will undergo efficient superradiant growth after $\psi_{211}$ saturates and dominate black hole spin down. This process is then followed by the superradiant growth of $\psi_{322}$ and $\varphi_{322}$. If $\gamma_{\varphi_{322}}>\gamma_{\psi_{322}}$ and $\tilde{a}^{\varphi_{211}}_{\rm crit}>\tilde{a}^{\psi_{211}}_{\rm crit},\tilde{a}^{\varphi_{322}}_{\rm crit}<\tilde{a}^{\psi_{211}}_{\rm crit}$ (region III), $\varphi_{211}$ will not grow, while $\varphi_{322}$ will grow efficiently after $\psi_{211}$ saturates, followed by the superradiant growth of $\psi_{322}$. A similar analysis can be extended in region IV, where higher modes of $\varphi$, such as $\varphi_{433}$, should also be taken into account. A representative example of each type is shown in Fig. 3. These examples are obtained by solving Eq. (10) together with $\dot{\epsilon}_{i}=\gamma_{i}\epsilon_{i}$, assuming the initial conditions of $\epsilon_{i}=10^{-10}$ when the black hole spin $\tilde{a}_{0}=0.99$. For simple, we treated the black hole mass (and hence $\alpha$) to be constant, given the fact that the black hole mass typically varies by a small amount during superradiant evolution. We find that, in region I and region IV, $\psi_{211}$ finally depletes after reaching the maximum occupation number, due to the development of $\varphi$ field.
@@ III Cubic Interactions
Assuming both fields are initially in vacuum states with small quantum fluctuations, the superradiance process can be investigated perturbatively: At the very beginning, interactions between the field is suppressed due to the low occupation numbers, and the superradiance process is the same as it in the case of free field. As superradiance continues, the occupation number in the fastest growing mode, i.e., ${\bar{\psi}}_{211}$, increases, resulting in a notable interaction between the two fields. Depending on the parameters, the interaction may lead to different processes: The $\psi$ field may lose energy to the black hole via the bound state of $\varphi$. It may also lose energy to the spatial infinity via the unbound states of $\varphi$ and $\psi$. The effects on the superradiant growth of such processes can be investigated by their corrections on the eigenfrequncies, which will be discussed as follows.
III.1 Bound state interactions
We shall start with interactions between $\varphi$ and the bound states of $\psi$. It is convenient to described the bound states with $\bar{\psi}$} defined in the non-relativistic ansatz (6), and work in the weak gravity limit. By substituting ansatz (6) and expanding terms in powers of $\alpha$, Eqs. (13) and (14) can be written into the form of
$\left(i\partial_{t}-\hat{H}\right)\left({\bar{\psi}}e^{-i\mu t}+{\bar{\psi}}^{*}e^{i\mu t}\right)=-2\bar{\lambda}\left({\bar{\psi}}e^{-i\mu t}+{\bar{\psi}}^{*}e^{i\mu t}\right)\varphi\,$, (15)
$\left(\Box-\nu^{2}\right)\varphi=\bar{\lambda}\left({\bar{\psi}}^{2}e^{-2i\mu t}+2{\bar{\psi}}{\bar{\psi}}^{*}+{{\bar{\psi}}}^{*2}e^{2i\mu t}\right),$
where $\bar{\lambda}\equiv\lambda/2\mu$ is the reduced coupling strength, and $\hat{H}\simeq-\tfrac{\nabla^{2}}{2\mu}-\tfrac{\alpha}{r}$ to the leading order of $\alpha$ (see App. A for the detailed derivation). Here we do not take the non-relativistic ansatz for $\varphi$, as its relativistic states could also contribute in this process.
To compute the corrections on the eigenfrequency, we take one of the $\bar{\psi}$} in the source terms as background ${\bar{\psi}}_{0}$. In our case, the background is mostly in the ${\bar{\psi}}_{211}$ mode, i.e., ${\bar{\psi}}_{0}=\sqrt{N_{0}}u_{211}(\mathbf{r})e^{-i\xi_{211}t}$. While ${\bar{\psi}}_{211}$ should in principle grow with time, given the fact that ${\rm Im}\xi_{211}\ll{\rm Re}\xi_{211}$, one can neglect the imaginary part of $\xi_{211}$ and treat ${\bar{\psi}}_{0}$ to be stationary in our perturbative calculation for simplicity [8]. It is convenient to use Feynman-like diagrams to represent the channels involved in Eqs. (LABEL:eq:eomm). In particular, $\bar{\psi}$} on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the l.h.s. of the vertex, and vice versa; and $\bar{\psi}$$^{*}$ on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the r.h.s. of the vertex, and vice versa. We are interested in the interaction corrections on the fastest growing mode ${\bar{\psi}}_{211}$. Given Eqs. (LABEL:eq:eomm), there are three possible channels, which are shown in Fig. 4.
@@ III.1 Bound state interactions
$\psi_{211}$$\psi_{211}$$\psi_{211}$$\psi_{211}$
Figure 4: Possible channels involved in Eqs. (LABEL:eq:eomm). The left diagram corresponds to the s-channel described by Eqs. (16) and (17), while the middle and right diagrams correspond to the t- and u-channels described by Eqs. (32).
@@
For the s-channel, we have
$\left(i\partial_{t}-\hat{H}\right){\bar{\psi}}=-2\bar{\lambda}{\bar{\psi}}_{0}^{*}\varphi e^{2i\mu t}\,$, (16)
$\left(\Box-\nu^{2}\right)\varphi=\bar{\lambda}{\bar{\psi}}{\bar{\psi}}_{0}e^{-2i\mu t}\,.$ (17)
To investigate the interaction corrections on $\omega_{211}$, we shall consider
${\bar{\psi}}={\bar{\psi}}^{(0)}+{\bar{\psi}}^{(1)}+\cdots$ (18)
with ${\bar{\psi}}^{(0)}=u_{211}e^{-i\xi_{211}t}$ and solve Eqs. (16) and (17) perturbatively. Here the superscription denotes the order of the perturbation calculation.
We shall start with Eq. (17). To the leading order in $\bar{\lambda}$, the r.h.s. of Eq. (17) is given by $\bar{\lambda}{\bar{\psi}}^{(0)}{\bar{\psi}}_{0}e^{-2i\mu t}$. Then by considering the ansatz
$\varphi(t,\mathbf{r})=\Phi(\mathbf{r})\,e^{-i\sigma_{0}t}\,$, (19)
Eq. (17) leads to $\sigma_{0}=2\omega_{211}$ and
$\left(\hat{M}_{\sigma=\sigma_{0}}-\nu^{2}\right)\Phi=\tilde{s}\left(\mathbf{r}\right)\,$, (20)
with $\hat{M}_{\sigma}$ can be obtained by replacing $\partial_{t}$ in $\Box$ with $i\sigma$ and $\tilde{s}\left(\mathbf{r}\right)=\bar{\lambda}\sqrt{N_{0}}u_{211}^{2}$. The above equation can be solved with the Green’s function method, [^3] Technically, the integral should be performed with the tortoise coordinates $\mathbf{r}_{*}$. See App. B for the precise formula. Nevertheless, we shall not distinguish $\mathbf{r}_{*}$ with $\mathbf{r}$ here, as $\tilde{s}(\mathbf{r})$ only has support around $r\sim GM/\alpha^{2}$ where $\mathbf{r}_{*}\simeq\mathbf{r}$.
$\Phi(\mathbf{r})=\int{\rm d}^{3}\mathbf{r}\,G_{\sigma=\sigma_{0}}(\mathbf{r},\mathbf{r}^{\prime})\,\tilde{s}\left(\mathbf{r}^{\prime}\right)\,.$ (21)
In App. B, we show that the above integral can be written as
$\Phi(\mathbf{r})=\sum_{n,\ell,m}^{\infty}b_{n\ell m}v_{n\ell m}(\mathbf{r})+\sum_{\ell,m}\int_{0}^{\infty}\mathrm{d}k\,b_{\ell m}(k)v_{k\ell m}(\mathbf{r})\,$, (22)
where $v_{n\ell m}$ and $v_{k\ell m}$ satisfying
$\left(\hat{M}_{\sigma=\sigma_{n\ell m}}-\nu^{2}\right)v_{n\ell m}=0\,,\quad\left(\hat{M}_{\sigma=\sigma_{k\ell m}}-\nu^{2}\right)v_{k\ell m}=0\,$, (23)
are the bound and scattering state wavefunctions of the free $\varphi$ field. Here the bound states has ${\rm Re}\sigma_{n\ell m}<\nu$ and are discrete in spectrum, while the scattering states has ${\rm Re}\sigma_{k\ell m}\geq\nu$ and are continuous in spectrum. [^4] Also see Ref. [64] for a recent discussion on the complete set of mode functions. The precise formulae for the coefficients $b_{n\ell m}$ and $b(k)_{\ell m}$ are derived in App. B. In practice, when we compute the numerical value of the coefficients, we approximate the wavefunctions $v_{n\ell m}$ and $v_{k\ell m}$ with their analytical expressions in the weak gravity limit (discussed in App. A), since $\tilde{s}(\mathbf{r})$ only has support far from the black hole horizon. It turns out that the coefficients are non-trivial only if $\ell=m=2$ due to the angular dependence, and are given by
$b_{n}$ $\simeq\frac{\bar{\lambda}\sqrt{N_{0}}}{\sigma_{n\ell m}^{2}-\sigma_{0}^{2}}\int u_{0}^{2}v_{n22}^{*}\,\mathrm{d}^{3}\mathbf{r}$ (24)
$b(k)$ $\simeq\frac{1}{2\pi}\frac{\bar{\lambda}\sqrt{N_{0}}}{\sigma_{k\ell m}^{2}-\sigma_{0}^{2}+}\int u_{0}^{2}v_{k22}^{*}\,\mathrm{d}^{3}\mathbf{r}$
where we have omitted the subscriptions $\ell m$} for short.
With the solution of $\varphi$, we now proceed to Eq. (16). Without lose of generality, we write
${\bar{\psi}}^{(1)}=u_{211}(\mathbf{r})e^{-i\delta\omega_{211}^{s}t}+\sum_{n\ell m\neq 211}a_{n\ell m}\,u_{n\ell m}(\mathbf{r})e^{-i\omega_{211}t}\,$, (25)
where $u_{n\ell m}$ denotes the bound states of $\bar{\psi}$}, and the summation excludes $u_{211}$ as which can be absorbed into ${\bar{\psi}}^{(0)}$. Substituting Eq. (25) into Eq. (16), and using the orthogonality of $u_{n\ell m}$, we can find $a_{n\ell m}=0$ and the frequency correction from the s-channel $\delta\omega_{211}^{s}$. Specifically, the correction can be written as $\delta\omega_{211}^{s}=\delta\omega_{211}^{sB}+\delta\omega_{211}^{sE}$ with
@@
${\rm Im}\delta\omega_{211}^{sB}$ (27)
$=$ $2\bar{\lambda}^{2}N_{0}\sum_{n\geq 3}{\rm Im}\left(\frac{1}{k_{0}^{2}+\kappa_{n}^{2}}\right)\left|\int u_{0}^{2}v_{n22}^{*}\mathrm{d}^{3}\mathbf{r}\right|^{2}$
$\simeq$ $2\bar{\lambda}^{2}N_{0}\frac{2q^{2}\alpha^{3}\mu}{(q^{2}-4)^{2}}\sum_{n\geq 3}\frac{{\rm Im}\sigma_{n}}{\nu}{\rm I}_{n}^{B}(q)\,$,
which indicates that the superradiant growth rate of ${\bar{\psi}}_{211}$ is suppressed if it interacts with a decaying ${\bar{\varphi}}_{n22}$, as ${\bar{\varphi}}_{n22}$ may lose energy to the black hole in the s-channel, and vice versa. In the last line of Eq. (27), we considered that $\omega_{0}\simeq\mu$ and ${\rm Re}\sigma_{n}\simeq q\mu$, and have defined
@@
which is a dimensionless factor depending only on the mass ratio $q$, with its values shown in Fig. 5. For the corrections from unbound states, we find that $\delta\omega_{211}^{sE}$ has a non-zero imaginary part only if $k_{0}$ is real due to the optical theorem. Specially, by performing the integral over $k$ with residue theorem, one has
${\rm Im}\delta\omega_{211}^{sE}\simeq\left\{\begin{array}[]{ll}-\bar{\lambda}^{2}N_{0}\frac{\mu\alpha^{2}}{2\sqrt{4-q^{2}}}{\rm I}^{E}_{k_{0}}(q,\alpha),\quad&2\omega_{0}>\nu\\ 0,\quad&2\omega_{0}<\nu\end{array}\right$.\,, (29)